OpenStudy (openstudier):

In the land of Alpha, (1/4)X(32)=5. What is (1/3)X(10)? A. 1 B. 2 C. 3 D. 4

OpenStudy (phi):

I assume this means 32 is not base 10. We need to find the base. 32 = 3*x +2 whatever the base, it is more than 5. multiply both sides by 4 3x+2 = 4*5=20(base 10) solve for x 3x= 18 x=6 base 6 works. now (1/3) * 10(base 6) is 1/3 * 6 =2

OpenStudy (openstudier):

I don't understand how you got the formula to find the base.

OpenStudy (openstudier):

Why do you solve for 32?

OpenStudy (openstudier):

Can you solve for the 1/4 or only just 32 to find the base?

OpenStudy (phi):

Example: a number abc in base 10 represents a*10^2 + b*10 + c abc in base x represents a*x^2 + b*x +c if we have just 2 digits: ab, this in base x stands for a*x + b

OpenStudy (openstudier):

I understand that.

OpenStudy (phi):

10 in base x represents the number x all numbers less than x will use one symbol. since they use 5, the base is bigger than 5 (and 4) so it safe to say 4 is the same in base x or in base 10 (it is four) and ditto for 5 we can still do algebra, so 4*1/4 * (3x+2) = 4*5 4/4 is one (even in base x) 4*5 is 20 in base 10

OpenStudy (openstudier):

So how do we solve that for the second part?

OpenStudy (phi):

I changed everything to base 10 to do the arithmetic. then change the answer back to base 6

OpenStudy (openstudier):

Ok. I understand. Thank you!